OpenStudy (anonymous):

solve the system of equations by graphing. Then classify the system. 3x + y = 9 4x +5y =1

OpenStudy (anonymous):

y = 9-3x y = (1-4x)/5

OpenStudy (anonymous):

9-3x = (1-4x)/5 45-15x=1-4x 44=11x x=4 y=-3

OpenStudy (anonymous):

What do they mean by classify it?